OpenStudy (okdutchman7):

Please help. WILL MEDAL! Mudog can do 8 jobs in 3 days, and Jimmy can do 5 jobs in 2 days. 39 jobs need to be done. Mudog works 3 days and then Jimmy joins in. How many days will both of them have to work together to complete the 39 jobs? @mathmale @TheSmartOne @sleepyjess @Directrix @mathstudent55 @Mehek14 @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

Is a formula given?

OpenStudy (okdutchman7):

I know that I have to find the rates of the the workers.

OpenStudy (okdutchman7):

It based on Distance=Rate x time

OpenStudy (okdutchman7):

\[Rate _{M} = \frac{ 8 jobs }{ 3 days}\]

OpenStudy (okdutchman7):

\[Rate _{J}=\frac{ 5 jobs }{ 2 days }\]

OpenStudy (okdutchman7):

\[Rate _{both} = \frac{ 31 jobs }{ 6 days }\]